OpenStudy (anonymous):

Please Help! Square root of 18a^5b^2

OpenStudy (ranga):

Factor out perfect squares: 18 = 9 * 2 (9 is a perfect square) a^5 = a^4 * a b^2 is already a perfect square. sqrt(18a^5b^2) = sqrt(9 * 2 * a^4 * a * b^2) = ?

OpenStudy (ranga):

sqrt(9) = 3 sqrt(a^4) = a^2 sqrt(b^2) = b The rest will remain inside the square root sqrt(9 * 2 * a^4 * a * b^2) = 3a^2b * sqrt(2a)

OpenStudy (ranga):

You can also write square root as exponent 1/2 and use the exponent rules to solve the problem.